OpenStudy (anonymous):

find a function f that satisfies the condition f''(x) = 2, f'(2) = 5 , f(2) = 10 this is antiderivatives and indefinite integrals (ie. applications of it)

OpenStudy (dumbcow):

f'(x) = 2x +C f'(2) = 5 ==> 2(2) +C = 5 ==> C = 1 f'(x) = 2x + 1 f(x) = x^2 + x + C f(2)=10 ==> 2^2 +2 +C = 10 ==> C=4 f(x) = x^2 + x +4

OpenStudy (anonymous):

thank you! i understand